Feather Friends, Inc., distributes a high-quality wooden birdhouse that sells for $120 per unit. Variable expenses are $60.00 per unit, and fixed expenses total $200,000 per year. Its operating results for last year were as follows:

Sales $ 3,240,000
Variable expenses 1,620,000
Contribution margin 1,620,000
Fixed expenses 200,000
Net operating income $ 1,420,000

Required:

Answer each question independently based on the original data:

1. What is the product's CM ratio?

2. Use the CM ratio to determine the break-even point in dollar sales.

3. If this year's sales increase by $44,000 and fixed expenses do not change, how much will net operating income increase?

4-a. What is the degree of operating leverage based on last year's sales?

4-b. Assume the president expects this year's sales to increase by 18%. Using the degree of operating leverage from last year, what percentage increase in net operating income will the company realize this year?

5. The sales manager is convinced that a 12% reduction in the selling price, combined with a $61,000 increase in advertising, would increase this year's unit sales by 25%.

a. If the sales manager is right, what would be this year's net operating income if his ideas are implemented?

b. If the sales manager's ideas are implemented, how much will net operating income increase or decrease over last year?

6. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $2.20 per unit. He thinks that this move, combined with some increase in advertising, would increase this year's sales by 25%. How much could the president increase this year's advertising expense and still earn the same $1,420,000 net operating income as last year?
Answer #1

1. What is the product's CM ratio?

CM ratio = Contribution / Sale *100

        = $1620000/$3240000 * 100 = 50%

2. Use the CM ratio to determine the break-even point in dollar sales.

Break even point = Fixed cost/CM ratio

         = $200000/50% = $400000

3. If this year's sales increase by $44,000 and fixed expenses do not change, how much will net operating income increase?

Sale Increase = $44000

Operating income increase = $44000 * 50% = $22000

4-a. What is the degree of operating leverage based on last year's sales?

Contribution/Operating Income

$1620000/$1420000 = 1.14

4-b. Assume the president expects this year's sales to increase by 18%. Using the degree of operating leverage from last year, what percentage increase in net operating income will the company realize this year?

Percentage Increase in net operating income = Percent of increase in sale * DOL

           = 18 * 1.14 = 20.52%

5a. If the sales manager is right, what would be this year's net operating income if his ideas are implemented?

New Selling price = $120 * 88% = $105.60

New sale in units = 27000 * 125% = 33750 units

Variable cost per unit = $60

Net operating income = Units sold * (SP – VC) – Fixed cost – Advertising cost

              = 33750 * ($105.60 - $60) - $200000 - $61000 = $1278000

5b. If the sales manager's ideas are implemented, how much will net operating income increase or decrease over last year?

Net operating income decrease = $1420000 - $1278000 = $142000

6. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $2.20 per unit. He thinks that this move, combined with some increase in advertising, would increase this year's sales by 25%. How much could the president increase this year's advertising expense and still earn the same $1,420,000 net operating income as last year?

New VC pu = $60 + $2.20 = $62.20

Net operating income = Units sold * (SP – VC) – Fixed cost – Advertising cost

$1420000 = 33750 * ($120 - $62.20) - $200000 – Advertising cost

$1420000 = $1950750 - $200000 – Advertising cost

Advertising cost = $330750
Answer #2
Req 1Product CM Ratio50.00%









Note 1:




CM Ratio = Contribution margin / sales




Contribution margin           1,620,000



Divided by sales           3,240,000



CM Ratio50.00%








Req 2Break Even points in sales dollar              400,000









Note 2:




Break Even points in sales dollar = Fixed expense/CM Ratio




Fixed expenses              200,000



Divided by CM Ratio50.00%



Break Even points in sales dollar              400,000








Req 3Net Operating Income Increase by           2,640,000









Note 3:




Effect on income = change in sales * CM Ratio




Increase in sales           5,280,000



Multiply by CM Ratio50%



Net Operating Income Increase by           2,640,000








Req 4-aDegree of Operating Leverage                       1.14



Note 4:




Operating Leverage = Contribution margin/ Net Income




Contribution Margin           1,620,000



Divided by Net Income           1,420,000



Degree of Operating Leverage                       1.14








Req 4-bNet Operating Income(NOI) increase by20.52%



Note 5:




Change in NOI = %age change in sales * operating leverage




%age change in sales18%



Multiply by operating leverage                       1.14



Net Operating Income(NOI) increase by20.52%








Req 5-aContribution Income   statement


 Last Year Proposed


                 27,000 Units               33,750 Units 


Total Per Unit Total Per Unit 

Sales           3,240,000                      120.00        3,564,000              105.60

Less: Variable expense           1,620,000                         60.00        2,025,000                 60.00

Contribution margin           1,620,000                         60.00        1,539,000                 45.60

Fixed expenses              200,000
            261,000

Net Income           1,420,000
        1,278,000






Req 5-bNet Operating Income Decrease by            (142,000)








Req 6The amount by which advertising can be increased              330,750









Note6:




Contribution margin per unit (Last year)                   60.00



Less: increase in sales comm                     (2.20)



Prposed contribution margin                   57.80



Multiplied by proposed sales after increase                 33,750



Proposed total contribution margin           1,950,750



Less: Fixed expenses            (200,000)



Less: Target Income        (1,420,000)



The amount by which advertising can be increased              330,750



source: managerial accounting
answered by: anonymous